### v1

@@ Abstract
This paper presents a significant improvement on the previous conference paper known as $DefSent$. The prior study seeks to improve sentence embeddings of language models by projecting definition sentences into the vector space of dictionary entries. We discover that this approach is not fully explored due to the methodological limitation of using word embeddings of language models to represent dictionary entries. This leads to two hindrances. First, dictionary entries are constrained by the single-word vocabulary, and thus cannot be fully exploited. Second, semantic representations of language models are known to be anisotropic, but pre-processing word embeddings for $DefSent$ is not allowed because its weight is frozen during training and tied to the prediction layer. In this paper, we propose a novel method to progressively build entry embeddings not subject to the limitations. As a result, definition sentences can be projected into a quasi-isotropic or isotropic vector space of unlimited dictionary entries, so that sentence embeddings of noticeably better quality are attainable. We abbreviate our approach as $DefSent+$ (a plus version of $DefSent$), involving the following strengths: 1) the task performance on measuring sentence similarities is significantly improved compared to $DefSent$; 2) when $DefSent+$ is used to further train data-augmented models like SIMCSE and SNCSE, state-of-the-art performance on measuring sentence similarities can be achieved among the approaches without using manually labeled datasets; 3) $DefSent+$ is also competitive in feature-based transfer for NLP downstream tasks.
@@ 1.1 Background
As token representations output from raw pre-trained BERT and RoBERTa suffer from anisotropic vector spaces (Gao et al., 2019; Ethayarajh, 2019), sentence embeddings encoded by raw pre-trained models via pooling strategies tend to carry similar semantic features for all sentences, and consequently poorly distinguish semantic similarities of sentence pairs, which is experimentally shown in Reimers and Gurevych (2019)’s work. To improve the sentence embeddings, there are three categories of finetuning raw pre-trained models:
@@ 1.2 Mechanism, motivation & research Goal
Given a dictionary dataset containing entries $e^{{}^{\prime}}$, for any entry-sentence pair ($e$, $s$) in the dataset, modeling $P_{\theta}(e|s)$ is formulated by Equation 1, where $h_{\theta}(s)$ is the representation output from a model to approximate $entry\_embed(e)$ and $\theta$ is trainable model parameters. $h_{\theta}(s)$ can also be interpreted as the representation projected into the vector space of dictionary entries.
$\operatorname*{arg\,max}\limits_{\theta}P_{\theta}(e|s)=\operatorname*{arg\,min}\limits_{\theta}-log\frac{exp(entry\_embed(e)\cdot h_{\theta}(s))}{\sum\limits_{e^{{}^{\prime}}}exp(entry\_embed(e^{{}^{\prime}})\cdot h_{\theta}(s))}$ (1)
@@ 1.4 Contributions
    Recent studies (Seonwoo et al., 2023) explore approaches to further training data-augmented models like SIMCSE (Gao et al., 2021) and SNCSE (Wang and Dou, 2023). We also show that when $DefSent+$ is used to further train those models, state-of-the-art performance can be achieved on STS tasks, among the approaches without using manually labeled datasets.
@@ 4 Unsupervised semantic textual similarity tasks
STS tasks without using any training data are used to test the capability of our sentence embeddings for the major application. The tasks cover STS 2012-2016 (Agirre et al., 2012, 2013, 2014, 2015, 2016), STS benchmark (Cer et al., 2017), and SICK-Relatedness (Marelli et al., 2014). The metric is Spearman’s rank correlation $\rho$ 100 between cosine similarities of sentence embeddings and human ratings.
@@
The nomenclature of $DefSent+$ encoders is $DefSent+^{PST-(pooling,type)-index}_{model\_name}$; e.g., $DefSent+^{1st-(CLS,AMP)-a}_{BERT-base}$ means that the experimental results are achieved by a BERT-base-uncased model trained from 1st Training of PST using Dataset $a$, and the encoding combination used during training is (CLS, $AMP$). After the final step of PST, for instance, the “ICA” in the lower brackets of $DefSent+^{3rd-(CLS,AMP)-a}_{BERT-base(ICA)}$ indicate that results are from training with the ICA-transformed vector space. If a “*” is attached to a encoder, the encoder will be released by us and also will be tested in Section 5 for the capability of feature-based transfer.
@@ 4.1 DefSent+ for raw pre-trained models
Table 3: The best results out of 10 runs are reported for $DefSent+$ encoders trained from data-augmented models: $SIMCSE$ (unsupervised) (Gao et al., 2021) and $SNCSE$ (Wang and Dou, 2023). The results of data-augmented models are based on their released encoders. For $SIMCSE$, we also attempt Prompt pooling to encode sentence embeddings, and report the results accordingly if average scores overtake their previous ones.
@@
Figure 8: The vector spaces of entry embeddings used to train the encoders from the data-augmented models.
@@ 4.2 DefSent+ for data-augmented models
In this section, we apply $DefSent+$ to data-augmented models, and the experimental results of STS tasks are reported in Table 3. Two sets of data-augmented models are further trained by $DefSent+$: unsupervised $SIMCSE$ (Gao et al., 2021) (the classical one preceding other CSE series) and $SNCSE$ (Wang and Dou, 2023) (the latest published one). We directly summarize the optimal vector spaces of entry embeddings used for training in Figure 8.
@@
Compared to their raw pre-trained counterparts, both difference and invariability are observed. For those previously relying on ICA-transformed (BERT encoders) and quasi-isotropic (RoBERTa-base encoder) vector spaces for training, the optimal distribution in vector spaces turn into a gourd-like shape (a big sphere-like at bottom with a small one at top). There is no sign of correlation between the distributions on different semantic axis. Plus, if the weight matrix of entry embeddings is centered first and then used for training, same level performance on STS tasks in Table 3 can be achieved. For these two considerations, The gourd-like distribution can also be considered as quasi-isotropic. The invariable pattern is observed from the optimal vector space of entry embeddings used for training RoBERTa-large encoders. To achieve the highest performance on STS tasks, 2nd Training of PST is needed, where entry distributions resemble rotated ellipse with a sign of negative correlation. This is not a coincidence, and we will present in the next section that RoBERTa encoders trained using this type of space are also powerful on the capability of feature-based transfer.
@@
The performance increments as shown in Table 8 are significant. Except for $DefSent+^{2nd-(Mean,AC)-a}_{sncse-RoBERTa-large}$ that improves the previous result by 0.4% (82.17% - 81.77%), the rest improvement ranges from 1.05% (80.28% - 79.23%) to 3.51% (79.76% - 76.25%). To the best of our knowledge, as of writing, the average results of STS tasks are the state-of-the-art performance achieved by the encoders trained without using large manually labeled datasets. There is one variation regarding encoding combinations. For further training $SNCSE_{RoBERTa-large}$, the encoding combination (Mean, AC) of 2nd Training is different from the one (CLS, AC) of 1st Training, but considering the fact that the total PST steps for further training data-augmented models are at most two, there is no big issue of consuming computational resources for one additional attempt. It seems that there is no noticeable performance difference between $DefSent+^{1st-(CLS,AC)-a}_{sncse-RoBERTa-large}$ and $DefSent+^{2nd-(Mean,AC)-a}_{sncse-RoBERTa-large}$; however, the statistical result of $DefSent+^{1st-(CLS,AC)-a}_{sncse-RoBERTa-large}$ ($81.81\pm 0.45$ in terms of average) is not comparable to the one of $DefSent+^{2nd-(Mean,AC)-a}_{sncse-RoBERTa-large}$ (see Table 6 in Appendix B).
@@
Table 4: Accuracy (%) performance on SentEval benchmarks with the default configuration. The results of previous works are based on their released encoders, and the results of our work are based on the encoders to be released. For $SIMCSE$ encoders, we attempt Prompt pooling to encode sentence embeddings, and report the results accordingly if average scores overtake their previous ones. For $SNCSE$ encoders, we attempt CLS or Mean pooling to encode sentence embeddings, and report the results accordingly if average scores overtake their previous ones. The results of $DefSent$ are based on their best pooling strategies.
@@ 5 SentEval benchmarks
We evaluate the capability of feature-based transfer for 12 $DefSent+$ encoders on SentEval (Conneau and Kiela, 2018) benchmarks. All the encoders are tested with the default configuration [^9] https://github.com/facebookresearch/SentEval run on the same GPU machine. Although the capability of feature-based transfer is not related to the major application of our sentence embeddings, it is still meaningful to further verify whether sentence embeddings encoded by $DefSent+$ encoders carries useful semantics, also as a way to double-confirm if data-augmented encoders are improved by $DefSent+$ or not.
The results are reported in Table 4. As shown in the table, sentence embeddings encoded by $DefSent$ encoders are more powerful than the ones encoded by data-augmented encoders, but are incomparable in terms of the performance on STS tasks regarding measuring similarities of sentence pairs. One the other hand, sentence embeddings encoded by $DefSent+$ encoders not only outperform $DefSent$ significantly on STS tasks, but also are comparable to $DefSent$ on the capability of feature-based transfer for various NLP downstream tasks. Two $DefSent+$ RoBERTa-large encoders achieve 89.00%+ average results, which are so far the state-of-the-art performance among the encoders not trained with large manually labeled datasets. Data-augmented encoders further trained by $DefSent+$ consistently outperform their previous counterparts in terms of average results, especially highlighted by the improvement achieved by RoBERTa-base encoders. This double-confirms that data-augmented encoders are improved by $DefSent+$. For the representative MRPC task, 5 $DefSent+$ encoders achieve the results over 77.60%, which are noticeably higher than 75.20% (purely based on lexical overlap features, as a token of textual similarities for semantic similarities) (Liu et al., 2023).
@@ 6 Conclusion
Among the approaches without using large manually labeled datasets, $DefSent+$ encoders achieve state-of-the-art performance on both STS tasks and SentEval benchmarks. The performance on STS tasks is achieved by further training data-augmented RoBERTa-large encoders based on unsupervised contrastive learning. The performance on SentEval benchmarks is achieved by training RoBERTa-large and further training data-augmented RoBERTa-large encoders.
@@ 7 Discussion & future directions
It can be argued that digital dictionary resources might be not available in all countries. Still, dictionary resources are comparatively easy to obtain, as there is no manual effort greatly in need to make annotations or labels. Moreover, the English dictionary resources we use in this paper are commonly-used ones, which are publicly available and not particularly tailored for our use.
Our GPU devices used for training are not the most advanced ones, but still acceptable in terms of the biggest Dataset $a$ with 1-epoch gradient descent: 48 minutes (RTX 2080 Ti) for BERT-base-uncased, 49 minutes (TITAN RTX) for RoBERTa-base, and 75 minutes (GeForce RTX 3090) for BERT-large-uncased and RoBERTa-large.
We will apply $DefSent+$ to other natural languages. Also, another two future directions, not the focus of this paper, can be investigated as well: exploration on the combination of other dictionary resources (as stated above, we focus on the commonly-used ones in this paper); application to reverse dictionaries (as explained in Section 3.2, in this paper, our focus is to encode sentence embeddings after training, and thus the optimized parameters of the pooler are no longer useful).

### v4

@@ Abstract
This paper presents a significant improvement on the previous conference paper known as $DefSent$. The prior study seeks to improve sentence embeddings of language models by projecting definition sentences into the vector space of dictionary entries. We discover that this approach is not fully explored due to the methodological limitation of using word embeddings of language models to represent dictionary entries. This leads to two hindrances. First, dictionary entries are constrained by the single-word vocabulary, and thus cannot be fully exploited. Second, semantic representations of language models are known to be anisotropic, but pre-processing word embeddings for $DefSent$ is not allowed because its weight is frozen during training and tied to the prediction layer. In this paper, we propose a novel method to progressively build entry embeddings not subject to the limitations. As a result, definition sentences can be projected into a quasi-isotropic or isotropic vector space of unlimited dictionary entries, so that sentence embeddings of noticeably better quality are attainable. We abbreviate our approach as $DefSent+$ (a plus version of $DefSent$), involving the following strengths: 1) the task performance on measuring sentence similarities is significantly improved compared to $DefSent$; 2) when $DefSent+$ is used to further train data-augmented models like SIMCSE, SNCSE, and SynCSE, state-of-the-art performance on measuring sentence similarities can be achieved among the approaches without using manually labeled datasets; 3) $DefSent+$ is also competitive in feature-based transfer for NLP downstream tasks.
@@ 1.1 Background
As token representations output from raw pre-trained BERT and RoBERTa suffer from anisotropic vector spaces (Gao et al., 2019; Ethayarajh, 2019), sentence embeddings encoded by raw pre-trained models via pooling strategies tend to carry similar semantic features for all sentences, and consequently poorly distinguish semantic similarities of sentence pairs, which is experimentally shown in Reimers and Gurevych (2019)’s work. To improve the sentence embeddings, there are three categories of fine-tuning raw pre-trained models:
@@ 1.2 Mechanism, motivation & research Goal
Given a dictionary dataset containing entries $e^{{}^{\prime}}$, for any entry-sentence pair ($e$, $s$) in the dataset, modeling $logp_{\theta}(e|s)$ is formulated by Equation 1, where $h_{\theta}(s)$ is the representation output from a model to approximate $entry\_embed(e)$ and $\theta$ is trainable model parameters. $h_{\theta}(s)$ can also be interpreted as the representation projected into the vector space of dictionary entries.
$\operatorname*{arg\,max}\limits_{\theta}logp_{\theta}(e|s)=\operatorname*{arg\,min}\limits_{\theta}-log\frac{exp(entry\_embed(e)\cdot h_{\theta}(s))}{\sum\limits_{e^{{}^{\prime}}}exp(entry\_embed(e^{{}^{\prime}})\cdot h_{\theta}(s))}$ (1)
@@ 1.4 Contributions
    Recent studies (Seonwoo et al., 2023) explore approaches to further training data-augmented models like SIMCSE (Gao et al., 2021), SNCSE (Wang and Dou, 2023) and SynCSE (Zhang et al., 2023). We also show that when $DefSent+$ is used for further training, state-of-the-art performance can be achieved on STS tasks, among the approaches without using manually labeled datasets.
@@ 4 Unsupervised semantic textual similarity tasks
STS tasks without using any training data are used to test the capability of our sentence embeddings for the major application [^9] Following DefSent (Tsukagoshi et al., 2021), we use the test datasets under /data/downstream of SentEval (Conneau and Kiela, 2018).. The tasks cover STS 2012-2016 (Agirre et al., 2012, 2013, 2014, 2015, 2016), STS benchmark (Cer et al., 2017), and SICK-Relatedness (Marelli et al., 2014). The metric is Spearman’s rank correlation $\rho$ 100 between cosine similarities of sentence embeddings and human ratings.
@@
The nomenclature of $DefSent+$ encoders is $DefSent+^{PST-(pooling,type)-index}_{model\_name}$; e.g., $DefSent+^{1st-(CLS,AMP)-a}_{BERT-base}$ means that the experimental results are achieved by a BERT-base-uncased model trained from 1st Training of PST using Dataset $a$, and the encoding combination used during training is (CLS, $AMP$). After the final step of PST, for instance, the “ICA” in the lower brackets of $DefSent+^{3rd-(CLS,AMP)-a}_{BERT-base(ICA)}$ indicate that results are from training with the ICA-transformed vector space. If a “*” is attached to an encoder, the encoder will be released by us and also will be tested in Section 5 for the capability of feature-based transfer.
@@ 4.1 DefSent+ for raw pre-trained models
Table 3: The best results out of 10 runs are reported for $DefSent+$ encoders trained from data-augmented models: $SIMCSE$ (unsupervised) (Gao et al., 2021), $SNCSE$ (Wang and Dou, 2023), and $SynCSE$ (Zhang et al., 2023). The results of data-augmented models are based on their released encoders. For $SIMCSE$ and $SynCSE$, we also attempt Prompt pooling to encode sentence embeddings, and report the results accordingly if average scores overtake their previous ones.
@@
[…]
$SynCSE(partial)_{RoBERTa-base}$
77.73 82.10 78.90 84.53 82.44 82.49 80.43 81.23
$SynCSE(partial)_{RoBERTa-large}$
75.95 84.95 80.20 85.80 84.52 85.26 82.20 82.70
@@
[…]
*$DefSent+^{1st-(Mean,AC)-a}_{syncse-partial-RoBERTa-base}$
77.32 83.94 81.10 85.99 83.73 83.82 79.50 82.20
$DefSent+^{1st-(CLS,AC)-a}_{syncse-partial-RoBERTa-large}$
75.63 86.21 80.18 85.62 84.42 85.03 78.75 82.26
@@
[…]
*$DefSent+^{2nd-(CLS,AC)-a}_{syncse-partial-RoBERTa-large}$
76.47 87.28 82.26 87.65 85.98 86.20 80.87 83.82
Figure 8: The vector spaces of entry embeddings used to train the encoders of the data-augmented models.
@@ 4.2 DefSent+ for data-augmented models
In this section, we apply $DefSent+$ to data-augmented models, and the experimental results of STS tasks are reported in Table 3. Three sets of data-augmented models are further trained by $DefSent+$: unsupervised $SIMCSE$ (Gao et al., 2021) (the classical one preceding other CSE series), $SNCSE$ (Wang and Dou, 2023) and $SynCSE$ (Zhang et al., 2023) (two latest published ones). The entry embeddings used for training are visualized in Figure 8.
@@
Compared to their raw pre-trained counterparts, both difference and invariability are observed. For those previously relying on ICA-transformed (BERT encoders) and quasi-isotropic (RoBERTa-base encoder) vector spaces for training, the optimal distribution in vector spaces turn into a gourd-like shape (a big sphere-like at bottom with a small one at top). There is no sign of correlation between the distributions on different semantic axis. Plus, if the weight matrix of entry embeddings is centered first and then used for training, same level performance on STS tasks in Table 3 can be achieved. For these two considerations, The gourd-like distribution can also be considered as quasi-isotropic. The invariable pattern is observed from the optimal vector space of entry embeddings used for training RoBERTa-large encoders. To achieve the highest performance on STS tasks, 2nd Training of PST is needed, where entry distributions resemble rotated ellipse with a sign of negative correlation. The consistency proves that RoBERTa-large encoders prefer this type of embeddings, and we will present in the next section that RoBERTa-large encoders trained using this type of embeddings are also powerful on the capability of feature-based transfer.
@@
The performance increments as shown in Table 8 are significant. Except for $DefSent+^{2nd-(Mean,AC)-a}_{sncse-RoBERTa-large}$ that improves the previous result by 0.4% ($82.17\%-81.77\%$), the rest improvement ranges from 0.97% ($82.20\%-81.23\%$) to 3.51% ($79.76\%-76.25\%$). To the best of our knowledge, as of writing, in terms of $DefSent+^{2nd-(CLS,AC)-a}_{syncse-partial-RoBERTa-large}$, the average result of STS tasks (83.82%) is the state-of-the-art performance achieved by the encoders trained without using large manually labeled datasets. There is one variation regarding encoding combinations. For further training $SNCSE_{RoBERTa-large}$, the encoding combination (Mean, AC) of 2nd Training is different from the one (CLS, AC) of 1st Training, but considering the fact that the total PST steps for further training data-augmented models are at most two, there is no big issue of consuming computational resources for one additional attempt. It seems that there is no noticeable performance difference between $DefSent+^{1st-(CLS,AC)-a}_{sncse-RoBERTa-large}$ and $DefSent+^{2nd-(Mean,AC)-a}_{sncse-RoBERTa-large}$; however, the statistical result of $DefSent+^{1st-(CLS,AC)-a}_{sncse-RoBERTa-large}$ ($81.81\pm 0.45$ in terms of average) is not comparable to the one of $DefSent+^{2nd-(Mean,AC)-a}_{sncse-RoBERTa-large}$ (see Table 6 in Appendix B).
@@
Table 4: Accuracy (%) performance on SentEval benchmarks with the default configuration is reported. The results of previous works are based on their released encoders, and the results of our work are based on the encoders to be released. For $SIMCSE$ and $SynCSE$ encoders, we attempt Prompt pooling to encode sentence embeddings, and report the results accordingly if average scores overtake their previous ones. For $SNCSE$ encoders, we attempt CLS or Mean pooling to encode sentence embeddings, and report the results accordingly if average scores overtake their previous ones. The results of $DefSent$ are based on their best pooling strategies.
@@
$SynCSE(partial)_{RoBERTa-base}$
85.77 91.63 93.21 90.23 90.94 93.00 77.22 88.86
$SynCSE(partial)_{RoBERTa-large}$
88.70 92.80 95.02 91.05 93.63 95.20 77.04 90.49
[…]
*$DefSent+^{1st-(Mean,AC)-a}_{syncse-partial-RoBERTa-base}$
85.70 91.50 93.82 90.69 91.32 92.20 77.80 89.00
*$DefSent+^{2nd-(CLS,AC)-a}_{syncse-partial-RoBERTa-large}$
87.88 92.13 94.53 90.82 92.86 94.20 77.68 90.01
@@ 5 SentEval benchmarks
We evaluate the capability of feature-based transfer for 14 $DefSent+$ encoders on SentEval (Conneau and Kiela, 2018) benchmarks. All the encoders are tested with the default configuration [^10] https://github.com/facebookresearch/SentEval run on the same GPU machine. Although the capability of feature-based transfer is not related to the major application of our sentence embeddings, it is still meaningful to further verify whether sentence embeddings encoded by $DefSent+$ encoders carries useful semantics, also as a way to double-confirm if data-augmented encoders are improved by $DefSent+$ or not.
The results are reported in Table 4. As shown in the table, sentence embeddings encoded by $DefSent$ encoders are more powerful than the ones encoded by data-augmented encoders ($SynCSE(partial)_{RoBERTa-large}$ is the only exception), but are incomparable in terms of the performance on STS tasks regarding measuring similarities of sentence pairs. One the other hand, sentence embeddings encoded by $DefSent+$ encoders not only outperform $DefSent$ significantly on STS tasks, but also are comparable to $DefSent$ on the capability of feature-based transfer for various NLP downstream tasks. Four $DefSent+$ encoders achieve 89.00%+ average results. Except $DefSent+^{2nd-(CLS,AC)-a}_{syncse-partial-RoBERTa-large}$ is outperformed by $SynCSE(partial)_{RoBERTa-large}$ [^11] When CLS pooling is used to encode sentence embeddings, $DefSent+^{2nd-(CLS,AC)-a}_{syncse-partial-RoBERTa-large}$ outperforms $SynCSE(partial)_{RoBERTa-large}$ by 0.65% ($89.34\%-88.69\%$), data-augmented encoders further trained by $DefSent+$ consistently outperform their previous counterparts in terms of average results, especially highlighted by the improvement achieved by RoBERTa-base encoders. This double-confirms that data-augmented encoders are improved by $DefSent+$. For the representative MRPC task, seven $DefSent+$ encoders achieve the results over 77.60%, which are noticeably higher than 75.20% (purely based on lexical overlap features, as a token of textual similarities for semantic similarities) (Liu et al., 2023).
@@ 6 Conclusion
Among the approaches without using large manually labeled datasets, $DefSent+$ encoders achieve state-of-the-art performance on both STS tasks. The performance on STS tasks is achieved by further training data-augmented RoBERTa-large encoders based on unsupervised contrastive learning.
@@ 7 Discussion & future directions
Data Availability. It can be argued that digital dictionary resources might be not available in all countries. Still, dictionary resources are comparatively easy to obtain, as there is no manual effort greatly in need to make annotations or labels. Moreover, the English dictionary resources we use in this paper are commonly-used ones, which are publicly available and not particularly tailored for our use.
Data Leakage. We confirm from two perspectives that there is NO data leakage (in machine learning sense) with regard to $DefSent+$. First, it is obvious that our training data are completely independent from test data — the training dataset comprises definition sentences while the test dataset contains sentences from task datasets (STS or SentEval). They were originally collected for different use cases. This further verifies that the training of $DefSent+$ generalizes well to unseen sentences. Second, it is widely acknowledged that neural networks entail the attribute called blackbox. When it comes to embeddings, the names of dimensionalities $\mathbb{d}$ (i.e., random variables; e.g., 768 $\mathbb{d}$ in BERT-base-uncased model) are unknown to us. Therefore, it is unviable to pre-define or select the $\mathbb{d}$ that are useful for the predictions of target labels. Furthermore, in this paper, we input entire embeddings into classifiers for various regression tasks without cherry-picking the $\mathbb{d}$.
Training Consumption. Our GPU devices used for training are not the most advanced ones, but still acceptable in terms of the biggest Dataset $a$ with 1-epoch gradient descent: 48 minutes (RTX 2080 Ti) for BERT-base-uncased, 49 minutes (TITAN RTX) for RoBERTa-base, and 75 minutes (GeForce RTX 3090) for BERT-large-uncased and RoBERTa-large.
Future Works. We will apply $DefSent+$ to other natural languages. Also, another two future directions, not the focus of this paper, can be investigated as well: exploration on the combination of other dictionary resources (as stated above, we focus on the commonly-used ones in this paper); application to reverse dictionaries (as explained in Section 3.2, in this paper, our focus is to encode sentence embeddings after training, and thus the optimized parameters of the pooler are no longer useful).